# Monte Carlo Simulation Development and Implementation of the GiBUU Model for Neutrino Experiments

### v1

@@ IV Unaddressed Developments in Systematical Uncertainty Estimation for GiBUU
Like any model, the GiBUU framework is subject to systematic uncertainties that can impact the precision of its predictions. After converting GiBUU into a Monte Carlo, the next step is to assess these systematic uncertainties and develop the software framework necessary to propagate those uncertainties. This section aims to delve into the areas of systematic uncertainty estimation that need to be designed for GiBUU. The inherent uncertainties associated with nuclear physics inputs can introduce systematic biases in the simulation results. Understanding and quantifying these uncertainties is crucial for robust data analysis and extracting meaningful physics results.
@@
We observe various sources contribute to systematic uncertainties in the GiBUU model:
Nuclear Effects and Medium Modifications.
@@
  • Medium Modifications: GiBUU considers the modification of hadron properties within the nuclear medium. This feature includes hadron masses and width changes due to interactions with surrounding nucleons. A deeper exploration of the dynamics of these medium modifications is necessary to refine uncertainty estimates.
Resonance Contributions.
@@
Neutrino-Nucleon Cross Sections.
Accurate cross-sections are crucial for reducing systematic uncertainties, but challenges persist:
@@
These studies’ comparisons between GiBUU simulations and experimental data are crucial in assessing the systematic uncertainties and validating the model. We can identify some of the associated uncertainties by comparing simulated event observables to data from electron and hadron scattering experiments. However, comparisons to neutrino scattering data are also critical.
@@ V Results and Discussion
We performed studies within this work using LArBox comparing GiBUU and GENIE predictions regarding the momentum (Fig. 4) and the cosine of the angle (Fig. 5) of outgoing muons in charged-current muon neutrino interactions. These figures highlight the nuanced differences between the models when generating events using the BNB neutrino flux at the LArBox detector location and its geometry.
@@
We calculated the ratio of different neutrino interaction topologies, such as $\nu_{\mu}CC0P0\pi^{\pm 0}$ (no protons or pions are observed in the final state above an energy threshold, defined as 50MeV in kinetic energy for protons and pions), $\nu_{\mu}CC1P0\pi^{\pm 0}$ (one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CCNP0\pi^{\pm 0}$ (more than one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{\pm}$ (one charged pion, no neutral pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{0}$ (one neutral pion, no charged pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CCOther$ (the remaining charged current events that didn’t fall into the previous categories), with respect to the inclusive $\nu_{\mu}CC$ events. The particle energy thresholds chosen are realistic if considering LArTPC detectors and large enough to reduce the impact due to the particle identification efficiency at lower energies.
@@
Figure 4 shows the ratio between a given neutrino-argon topology and the inclusive sample with respect to the muon momentum for GiBUU at the top and GENIE at the bottom. The most noticeable difference is that the contributions of each topology are clearly different in each model. In GiBUU, the dominant contributions are $\nu_{\mu}CC1P0\pi^{\pm 0}$ and $\nu_{\mu}CC0P0\pi^{\pm 0}$, while in GENIE the main contributions are due to $\nu_{\mu}CCNP0\pi^{\pm 0}$ and $\nu_{\mu}CCOther$.
By investigating using different particle energy thresholds and looking at different particle multiplicity, it is observed that GENIE produces higher proton multiplicity than GiBUU; the differences can be explained by the different neutrino interaction compositions per model, particularly for MEC and resonant productions, but as well the nuclear medium modeling.
Figure 5 shows similar ratios but with respect to the cosine of the muon angle with respect to the neutrino direction. We observe certain suppression of events in the most forward muons that affect mostly the $\nu_{\mu}CC1P0\pi^{\pm 0}$ topology in GiBUU and the $\nu_{\mu}CCNP0\pi^{\pm 0}$ in GENIE. These distributions are of special importance to understanding neutrino data, and the different trends of each Monte Carlo offer much information.
The different compositions by Monte Carlo tell us about the potential background predictions we can find in experimental data, which are very different depending on the model choice. These different composition predictions would lead to important corrections in the neutrino energy reconstruction for an oscillation experiment and could lead to different answers.
@@
Figure 4: Ratios between each exclusive topology and the inclusive charged-current muon neutrino predictions regarding the muon momentum comparing GiBUU and GENIE in the LArBox geometry using the BNB flux.
@@
Figure 5: Ratios between each exclusive topology and the inclusive charged-current muon neutrino predictions regarding the muon angle comparing GiBUU and GENIE in the LArBox geometry using the BNB flux.
@@
In high-resolution detectors, such as LArTPC, the final particle content observed in the detector is critical for reconstructing neutrino energy. This approach, aimed at reducing model-dependent uncertainties compared to methods such as CCQE-like approaches, poses a significant challenge in identifying all hadrons emerging from the nucleus. A major hurdle is the quantification of invisible neutrons and the complex modeling of final proton content and their kinematics. Correcting for these observables, especially for nucleons, can introduce substantial uncertainties and biases in neutrino energy reconstruction. Accurate modeling of the number of nucleons knocked out during neutrino interactions is critical. Access to diverse modeling approaches is essential for addressing this challenge. In Figure 6, we present the average number of protons generated per neutrino interaction across varying neutrino energies. Solid lines represent predictions from the GiBUU model, while dashed lines correspond to GENIE predictions. This comparison underscores the nuanced differences between these models that will directly reflect when reconstructing neutrino energies at experiments such as SBN and DUNE.
@@
Figure 6: Average number of protons simulated above 50 MeV per neutrino at each neutrino energy, per interaction type. Solid lines represent predictions from GiBUU; dashed lines represent predictions from GENIE.
@@
Validation against experimental data, including charged-current and neutral-current interactions, will demonstrate the model’s capacity to reproduce a broad range of observables faithfully. Our study reinforces the importance of model diversity in refining the interpretation of neutrino scattering data, which is essential for pushing the boundaries of our understanding in this intricate field.
@@ VI Conclusion
In conclusion, our implementation of the GiBUU model as a Monte Carlo has yielded valuable insights into the complex dynamics of neutrino scattering processes. This Monte Carlo reproduces correctly the GiBUU model, and it is the first step in incorporating uncertainties to be propagated into experimental data analysis. Notably, a comparative study with GENIE reveals nuanced differences, emphasizing the necessity of employing multiple models to interpret neutrino scattering data comprehensively.
@@ VII Acknowledgments
We thank Prof. Ulrich Mosel and Prof. Kai Gallmeister, from Justus-Liebig-Universitat Giessen, Germany, for their indispensable guidance. Their expertise significantly enhanced our study of neutrino interactions.

### v2

@@ IV Ongoing Developments in Systematic Uncertainty Estimation for GiBUU
Like any model, the GiBUU framework is subject to systematic uncertainties that can impact the precision of its predictions. After converting GiBUU into a Monte Carlo, the next step is to assess these systematic uncertainties and develop the software framework necessary to propagate those uncertainties. This section aims to delve into the areas of systematic uncertainty estimation that need to be designed for GiBUU. The inherent uncertainties associated with nuclear physics inputs can introduce systematic biases in the simulation results. Understanding and quantifying these uncertainties is crucial for robust data analysis and extracting meaningful physics results. For a deeper discussion on the models we will refer to [16].
@@
We observe that various sources contribute to systematic uncertainties in the GiBUU model:
Nuclear Effects and Medium Modifications
@@
  • Medium Modifications: GiBUU considers modifying hadron properties within the nuclear medium. This feature includes hadron masses and width changes due to interactions with surrounding nucleons. A deeper exploration of the dynamics of these medium modifications is necessary to refine uncertainty estimates.
Resonance Contributions
@@
Neutrino-Nucleon Cross Sections
Accurate cross sections are crucial for reducing systematic uncertainties, but challenges persist:
@@
Comparisons between GiBUU simulations and experimental data are crucial for assessing systematic uncertainties and validating the model. We can identify some of the associated uncertainties by comparing simulated event observables to data from electron and hadron scattering experiments. However, comparisons to neutrino scattering data are also critical.
The development of the software infrastructure for GiBUU systematics and its evaluation is an ongoing effort that will be described in a future publication. Currently, this work uses data to evaluate uncertainties and generate different event libraries with these variations.
A package we are implementing, GiBUU-Systematics, will handle all systematics and will interface with tools used in neutrino experiments. The main difference with other current neutrino generator reweighting tools is that GiBUU-Systematics produces systematic variations simultaneously with the central value library without interfering with each other, storing them in a separate library. This allows experiments to generate both central values and systematic variations per neutrino event, making them immediately available to analyzers. Producing systematic variations beforehand, rather than reweighting afterward, is a novel approach we believe offers benefits without compromising physics accuracy.
To assess uncertainties within GiBUU-Systematics, we employ a three-step evaluation process, taking advantage that GiBUU is not tuned to neutrino data. First, we compare with hadron scattering data, which is sensitive to strong interactions and uses highly accurate beams. The second step involves comparing with electron scattering data, which is sensitive to electroweak interactions and offers similar beam accuracy. These first two steps help capture uncertainties that affect the final state of the interaction and some accessible electroweak parameters.
In the third step, we compare with neutrino scattering data to capture uncertainties not visible in hadron and electron scattering. Neutrino interactions in nuclei are sensitive to weak interactions, though particles produced are also subject to strong interactions, impacting the event’s final state. Since factorization of effects in neutrino data is challenging and may be impossible, and beam uncertainties are significant due to neutrino production mechanisms, this comparison step is the final one.
@@ V Results and Discussion
In this work, we performed studies using LArBox to compare GiBUU and GENIE predictions for the momentum (Fig. 5) and the cosine of the angle (Fig. 7) of outgoing muons in $\nu_{\mu}CC$ interactions. These figures highlight the nuanced differences between the models when generating events with the BNB neutrino flux at the LArBox detector location and geometry.
@@
We calculated the ratio of different neutrino interaction topologies relative to the inclusive $\nu_{\mu}CC$ events. These topologies include $\nu_{\mu}CC0P0\pi^{\pm/0}$ (no protons or pions are observed in the final state above an energy threshold, defined as 50 MeV in kinetic energy for protons and pions), $\nu_{\mu}CC1P0\pi^{\pm/0}$ (one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CCNP0\pi^{\pm/0}$ (more than one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{\pm}0\pi^{0}$ (one charged pion, no neutral pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{0}$ (one neutral pion, any number of charged pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CCOther$ (remaining charged current events not included in the previous exclusive topologies). The chosen particle energy thresholds are realistic if considering LArTPC detectors and large enough to reduce the impact due to the particle identification efficiency at lower energies.
@@
Figure 5 shows the ratio between a given neutrino-argon topology and the inclusive sample with respect to the muon momentum for GiBUU at the top and GENIE at the middle. The most noticeable difference is that the contributions of each topology are clearly different in each model. In GiBUU, the dominant contributions are $\nu_{\mu}CC1P0\pi^{\pm/0}$ and $\nu_{\mu}CC0P0\pi^{\pm/0}$, while in GENIE, the main contribution is $\nu_{\mu}CC1P0\pi^{\pm/0}$, with clear differences in the shape of the distributions across topologies between the models. The third plot in Figure 5 is the ratio of the fractional contributions of each topology to the $\nu_{\mu}CC$ between GiBUU and GENIE to offer more details on the fractional differences between these two models.
@@
Figure 5: Ratios of each exclusive topology to the inclusive charged-current muon neutrino prediction as a function of muon momentum, comparing GiBUU and GENIE in the LArBox geometry with the BNB flux. The top and middle plots show the ratios for GiBUU and GENIE, respectively. The bottom plot displays the double ratio, calculated as the ratio between the top and middle plots. Proton and pion kinetic energy thresholds are set at 50 MeV to define the topologies.
By investigating different particle energy thresholds and particle multiplicities, we observed that GENIE produces a higher proton multiplicity than GiBUU. For example, as shown in Figure 6, GENIE has a lower composition of events with no protons below the energy threshold in the $\nu_{\mu}CC0\pi^{\pm/0}$ topology. We also observe different neutrino interaction compositions in each model, particularly in MEC and quasi-elastic production. These differences, along with variations in nuclear medium modeling, may explain the discrepancy in the $\nu_{\mu}CC0P0\pi^{\pm/0}$ topology between the two models, which exhibited the largest difference in the double ratio from Figure 5.
[…]
Figure 7 shows similar ratios, but with respect to the cosine of the muon angle relative to the neutrino direction. We observe a suppression of events for the most forward muons, which predominantly affect the $\nu_{\mu}CC1P0\pi^{\pm/0}$ topology in GiBUU and the $\nu_{\mu}CC0P0\pi^{\pm/0}$ topology in GENIE. These distributions are particularly important to understanding neutrino data, and the different trends of each Monte Carlo offer much information. The third plot shows the double ratio of these quantities, similar to Figure 5.
@@
Figure 7: Ratios of each exclusive topology to the inclusive charged-current muon neutrino prediction as a function of muon angle, comparing GiBUU and GENIE in the LArBox geometry with the BNB flux. The top and middle plots show the ratios for GiBUU and GENIE, respectively. The bottom plot displays the double ratio, calculated as the ratio between the top and middle plots. Proton and pion kinetic energy thresholds are set at 50 MeV to define the topologies.
The different compositions by Monte Carlo tell us about the potential background predictions we can find in experimental data, which are very different depending on the model choice. These different composition predictions would also lead to important corrections in the neutrino energy reconstruction for an oscillation experiment and could produce to different outcomes.
@@
In high-resolution detectors, such as LArTPC, the final observed particle content in the detector is critical for reconstructing neutrino energy. This approach, which aims to reduce model-dependent uncertainties compared to methods like CCQE-like approaches, faces significant challenges in identifying all hadrons emerging from the nucleus. A major hurdle is quantifying invisible neutrons and accurately modeling the final proton content and their kinematics. Correcting these observables, especially for nucleons, can introduce substantial uncertainties and biases in neutrino energy reconstruction.
Accurate modeling of the number of nucleons knocked out during neutrino interactions is essential, and access to diverse modeling approaches is crucial for tackling this challenge. In Figure 8, we present the average number of protons generated per neutrino interaction across varying neutrino energies. Solid lines represent predictions from the GiBUU model, while dashed lines correspond to GENIE predictions. Generally, we observe that GENIE has a higher proton multiplicity than GiBUU, except for DIS above $\sim 1.4GeV$.
@@
Figure 8: Average number of simulated protons above 50 MeV per neutrino at each neutrino energy, categorized by interaction type. Solid lines represent predictions from GiBUU, while dashed lines represent predictions from GENIE.
Notably, observing final-state protons assists in reconstructing the neutrino energy. For particles not observed—such as neutrons and low-energy particles below the detection threshold—model-based corrections are applied for oscillation measurements. This is one of the most model-dependent steps in oscillation data analysis, with the modeling of final-state nucleons being a substantial challenge for generators. This comparison highlights the nuanced differences between these models, which will directly impact neutrino energy reconstruction in experiments like SBN and DUNE.
@@
Validation against experimental data, including charged-current and neutral-current interactions, will demonstrate the model’s capacity to faithfully reproduce a broad range of observables. Our study reinforces the importance of model diversity in refining the interpretation of neutrino scattering data, which is essential for pushing the boundaries of our understanding in this intricate field. The work developed is currently used by the SBND and ICARUS collaborations, and the framework will soon be integrated into LArSoft, enhancing accessibility across experiments. It can also be easily integrated into other experiments that do not use LArSoft, as long as they use GENIE for their event generator. For experiments that do not use GENIE, integration remains feasible by employing their event generator’s flux and geometry drivers.
@@ VI Conclusion
In conclusion, implementing the GiBUU model as a Monte Carlo has yielded valuable insights into the complex dynamics of neutrino scattering processes. This Monte Carlo reproduces correctly the GiBUU model, and it is the first step in incorporating uncertainties to be propagated into experimental data analysis. Notably, a comparative study with GENIE reveals nuanced differences, emphasizing the necessity of employing multiple models to interpret neutrino scattering data comprehensively. The significant differences observed between GiBUU and GENIE, particularly in expected topological event samples and proton multiplicity (Figure 56, and  7), highlight the importance of understanding these models to better assess uncertainties when reconstructing neutrino energy, especially for oscillation experiments. GiBUU offers a different approach to other generators with a unified nucleus description, bringing new opportunities to better explore our data and develop model uncertainties. The same methodology we used to incorporate GiBUU as an unweighted event generator could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations. Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB flux.
@@ VII Acknowledgments
We thank Prof. Ulrich Mosel and Prof. Kai Gallmeister from Justus-Liebig-Universitat Giessen, Germany, for their indispensable guidance. Their expertise significantly enhanced our study of neutrino interactions.